OpenStudy (liliakarina):

How can I transform polar coordenates or rectangular? Here's an example: (6,45°)

OpenStudy (amistre64):

6 tells you how far it is from the origin: 6 = sqrt(x^2 + y^2) the degree tells us the slope of the line that it is on thru the origin: tan(45) = 1

OpenStudy (amistre64):

y = x is the line that its on

OpenStudy (amistre64):

some knowledge of a 45degree triangle would be useful here

OpenStudy (amistre64):

but regardless; y = x, so sub it into the distance from the origin part 6 = sqrt(x^2 + y^2) 6 = sqrt(x^2 + (x)^2) 36 = x^2 + x^2 36 = 2x^2 18 = x^2 x = sqrt(18) = 3 sqrt(2) and since y = x, the leaves us with the rectangulars (3sqrt(2),3sqrt(2))

OpenStudy (liliakarina):

Thank you, so (4.24, 4.24) is okay?

OpenStudy (amistre64):

i believe so ;)

OpenStudy (liliakarina):

All right, thanks for your help